OpenStudy (anonymous):

6. Write the equation of a line in slope intercept form that passes through the points (-2, 7) and (3, 12)

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Nnesha

OpenStudy (some.random.cool.kid):

hello :}

OpenStudy (anonymous):

fan and metal to helper

OpenStudy (some.random.cool.kid):

did they give a graph if not thats ok.

OpenStudy (zale101):

Use the slope formula \(slope=\Large\frac{y_2-y_1}{x_2-x_1}\) x1,y1= (-2, 7) and x2,y2=(3, 12) plug to find the slope

OpenStudy (zale101):

can you do that?

OpenStudy (zale101):

We are trying to find the slope in order to construct the equation. We don't have the slope y=mx+b where m is the slope and b is the y-intercept.

OpenStudy (some.random.cool.kid):

also i believe you can use the y = mx + b method to solve. am i correct zale?

OpenStudy (some.random.cool.kid):

oh nvm you said it

OpenStudy (zale101):

Yes,but we don't have the slope yet. @highjumpcollect needs to find it :)

OpenStudy (some.random.cool.kid):

yeah ik

OpenStudy (anonymous):

ok hold on let me solve it rn

OpenStudy (zale101):

Take your time :)

OpenStudy (anonymous):

oh wait i tried that and got 1 but i thought that was wrong

OpenStudy (some.random.cool.kid):

im in no rush im doing my own math right now too :}

OpenStudy (anonymous):

i had a toutor today to help me with this and we left off on this problem so I'm kinda forgetting how to do this

OpenStudy (zale101):

\(slope=\Large\frac{y_2-y_1}{x_2-x_1}=\)\(\Large\frac{12-(7)}{3-(-2)}=\Large\frac{12-7}{3+2}=\)\(\Large\frac{5}{5}=1\)

OpenStudy (zale101):

The slope is therefore 1 :)

OpenStudy (anonymous):

exactly what i did and got ok whats the next step i got lost there

OpenStudy (zale101):

Now, let's plug this into the slope intercept form y=mx+b where m (slope) is 1 y=1x+b to solve for the y intercept which is b in this case, we plug in (x,y) to one of the given point

OpenStudy (zale101):

i'll choose (-2,7)

OpenStudy (anonymous):

ok

OpenStudy (zale101):

\(y=1x+b\) \((7)=1(-2)+b\) add 2 to both sides \(7+2=b\) \(\Large b=9\)

OpenStudy (zale101):

y=1x+9

OpenStudy (anonymous):

omg thank u so much that makes a lot of sense could u help me with a couple more problems?

OpenStudy (zale101):

Have you tried solving them? Use the same procedure i did above and you'll do well...

OpenStudy (anonymous):

yes i did but it made no sense...

OpenStudy (zale101):

let me see what you did :)

OpenStudy (anonymous):

the problem is just like this one but i got 0 as my slope.. this one i did with my touter and idk what he was saying

OpenStudy (anonymous):

this is the problem - Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4).

OpenStudy (anonymous):

i did the